# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v1

@@ Abstract
Bridging the microscopic/individual and the macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control tasks. Here, we propose a combined manifold and machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations. The proposed framework builds upon our previous works on the development of next-generation Equation-free algorithms for dealing with the “curse of dimensionality” when learning surrogate models for high-dimensional and multiscale systems. The proposed approach is a four-stage one, explicitly conserving the mass of the reconstructed dynamics in the high-dimensional space. In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation (KDE). In the second step, based on manifold learning, we construct a map from the macroscopic ambient space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distribution. The third step involves learning reduced-order surrogate models (ROMs) in the latent space using machine learning techniques, particularly Long Short-Term Memory (LSTMs) networks and Multivariate Autoregressive models (MVARs). Finally, we reconstruct the crowd dynamics in the high-dimensional space in terms of macroscopic density profiles. We demonstrate that the POD reconstruction of the density distribution via SVD conserves the mass. With this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation (PDE) for the evolution of the density distribution. For our illustrations, we use the Social Force Model (SFM) to generate data in a corridor with an obstacle, imposing periodic boundary conditions. The numerical results demonstrate high accuracy, robustness, and generalizability, thus allowing for fast and accurate modelling/simulation of crowd dynamics from agent-based simulations. Notably, linear MVAR models surpass their nonlinear counterparts, such as LSTMs, in predictive accuracy, while also offering significantly lower complexity and greater interpretability.
Keywords Crowd Dynamics, Multiscale modelling, Micro to Macro, Machine Learning, Numerical Analysis, Conservation laws, Reduced Order Models, Manifold Learning
@@ 1 Introduction
An important open challenge in the modelling of human crowd dynamics lies in bridging across the different modelling scales. Typically, models are formulated at three distinct levels: the microscopic (individual), mesoscopic (kinetic), and macroscopic (hydrodynamic) scales. Microscopic models describe the detailed behavior of individual pedestrians using agent-based approaches, such as the Social Force Model (SFM) [1] and its variants (see e.g. [2, 3]), cellular automata [4, 5, 6], etc. The mesoscopic scale serves as an intermediary between the microscopic and macroscopic approaches [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. At the macroscopic scale, modelling describes the emergent/collective behavior of crowds. For instance, in the seminal work of Hughes [17], the author models crowd dynamics by coupling a conservation law with an eikonal equation to represent how individuals choose paths based on a potential field. Moreover, in [18] the authors proposed nonlocal conservation laws for traffic and crowd dynamics, where the flux depends not only on local density but also on averaged information from the surrounding region. In [11] a framework for macroscopic crowd models was presented based on balance laws and kinetic theory, emphasizing how individual-level interactions give rise to complex collective behavior. More recently, in [16] are highlighted new trends such as multiscale coupling, behavioral heterogeneity, and nonlocal interaction kernels.
@@
A major objective, when high-fidelity microsocpic/individual/agent-based simulators and/or detailed experimental data are available, is to derive mesoscopic or macroscopic models that can be then used for rigorous numerical analysis, optimization and control of the emergent/collective dynamics, for instance, in the optimal design of spaces, the reduction of evacuation times, and the management of crowd flows [11, 19, 20]. In this context, kinetic models serve as an intermediate bridge. They are typically derived analytically using statistical mechanics techniques, starting from microscopic interaction rules. These kinetic formulations, in turn, facilitate the derivation of macroscopic (continuum) models through asymptotic methods [10, 16].
@@
The transition from high-dimensional agent-based representations to macroscopic or hydrodynamic descriptions is commonly achieved through closure techniques. These methods establish relationships between higher-order moments of the microscopic distributions and a limited, smaller set of lower-order moments that capture the observable emergent behavior [21, 22, 23]. However, such closures inherently rely on assumptions that introduce biases both for modelling and analyses tasks. Typically, these assumptions include infinitely large populations, homogeneous agents, and uniform and/or local interaction networks, but also unknown “behavioural” characteristics, which can significantly impact the accuracy and validity of collective-level analyses.
@@
Recently, data-driven computational approaches based on physics-free machine learning (ML) have been employed for the modelling, numerical analysis, and control of pedestrian, crowd, and traffic dynamics. In this context, a wide range of ML architectures has been explored, including deep neural networks (DNNs) [24], Long Short-Term Memory (LSTM) networks [25, 26] for learning local movement decisions, Generative Adversarial Networks (GANs) [27, 28], Convolutional Neural Networks (CNNs) [29, 30, 31], and reinforcement learning methods [32, 33]. For a comprehensive review of these and related ML approaches, see [34]. These approaches have proven effective in forecasting individual and group trajectories within the framework of time-series modelling and forecasting, without relying on physics-based formulations. However, at the same time, such approaches lack interpretability and, importantly, do not incorporate information of the global “spatio-temporal” emergent dynamics, such as density fields, that are essential in modelling crowd dynamics.
@@
An alternative data-driven paradigm that bypasses the need to explicitly learn surrogate models is the Equation-free multiscale framework [35]. By appropriately coupling on-demand microscopic simulations with coarse-graining (and fine-graining), this approach builds a coarse time-stepper for the evolution at the macroscopic scale, thus bypassing the need for analytically derived continuum equations [36, 37, 38]. While Equation-free framework facilitates powerful multiscale numerical analysis and control, it lacks an explicit, global macroscopic equation for the crowd/collective dynamics–that can be approximated, such as a mass-conserving PDE–limits interpretability, physical insight, and generalization for long-term simulations.
@@
To address the aforementioned challenges, a few recent studies have utilized physics-informed neural networks (PINNs) [39] including physics-informed neural operators (NOs) [40, 41] for solving the inverse problem, i.e. that of constructing from data, hybrid models in the form of “grey-box” PDE models [42, 43, 21, 23, 41]. Such an approach, combines ML with underlying knowledge of physical principles, such as conservation laws [44] and closures [45], to create hybrid models that can both predict the spatio-temporal dynamics and ensure physical consistency, even under data scarcity. In the context of crowd dynamics, PINNs have been employed to construct both gray-box hydrodynamic PDEs [46] as well as for discovering the physics/rules governing interactions and motion at the microscopic level [47]. However, while PINNs provide an effective way to incorporate physical constraints and knowledge into machine learning models, they can also introduce biases due to their reliance on predefined forms of PDEs. Moreover, a key physical property in crowd modeling—mass conservation—is typically not enforced explicitly, but rather imposed as a soft constraint through the loss function [48, 44, 49]. This indirect enforcement can limit the generalizability and physical reliability of PINN-based models, particularly in heterogeneous or unforeseen scenarios where the true dynamics deviate from the assumed PDE forms.
@@
For learning classes of PDEs that incorporate conservation laws—such as those found in crowd dynamics—the use of black-box deep neural networks (DNNs) and black-box neural operators (NOs) [50, 51, 49, 52, 53, 23] is inherently problematic. These models often lack physical consistency, particularly in preserving conservation laws, which are usually enforced only as soft constraints in the loss function. In addition, their training suffers from low approximation accuracy due to the curse of dimensionality, potentially leading to unphysical predictions and numerical instabilities during long-term simulations. Thus, to deal with that problem, RiemannONets were recently presented in [54], which are NOs based on DeepONets to solve Riemann problems encountered in compressible flows, thus enforcing the preservation of conservation laws. Towards solving the inverse problem, we have recently proposed GoRINNs [55]: Godunov-Riemann informed neural networks for learning hyperbolic PDEs that respect the conservation laws explicitly.
It is important to note, however, that despite the long-standing tradition of modeling crowd dynamics through (hydrodynamic) PDEs, such formulations inherently rely on an “infinite-size” assumption. In realistic settings involving crowds of moderate or finite size, this assumption often breaks down. Finite-size effects—such as spatial heterogeneity, individual interactions, and boundary influences—can dominate the dynamics, rendering continuum PDEs insufficient for accurate, quantitative predictions of real-world crowd behavior.
@@
Here, inspired by the Equation-free multiscale framework [35], and building on our previous recent works on what we call next-generation Equation-free algorithms based on both manifold learning and machine learning [56, 36, 57, 23, 58, 41], we present a computational framework, to learn the discrete evolution operator of the emergent dynamics in latent spaces from high-fidelity microscopic simulations of crowd dynamics. The proposed framework bridges the microscopic and macroscopic modelling scales, while explicitly preserving the underlying conservation law of mass in the ambient space. The methodology is deployed in four main stages. First, we map discrete microscopic distributions to continuous macroscopic representations (i.e., density fields). Second, we project the macroscopic fields into a low-dimensional latent space using Proper Orthogonal Decomposition (POD). Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM networks. These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’ theorem [59] for reconstructing the phase-space structure (here in the latent space). The concept of the approach (i.e. that using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [60, 61] (see also results and discussion in [61, 56, 62, 63, 57]). In crowd dynamics, behavioural characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Taken’s theorem, to reconstruct the effective phase space containing these (unobservable) latent dynamics. Finally, we lift the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis. Crucially, this configuration ensures explicit mass conservation—a key physical constraint for any effective continuum description. For our numerical experiments, we employ the Social Force Model (SFM) [64] to simulate pedestrian flow within a rectangular corridor containing an obstacle, under periodic boundary conditions. The performance and generalization ability of the proposed framework are assessed on unseen initial conditions.
@@ 2.1 Problem Statement
We utilize a detailed simulator that models the dynamics of a human crowd, consisting of $N$ pedestrians moving within a domain $\Omega\subset\mathbb{R}^{2}$. The overall state of the pedestrian $i\in\{1,\ldots,N\}$ is characterized by its position $\mathbf{x}_{i}=\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ and velocity $\mathbf{v}_{i}=\mathbf{v}_{i}(t)=(v_{ix}(t),v_{iy}(t))$, and also a set of “behavioral” variables $\bm{u}_{i}=\bm{u}_{i}(t)\in\mathbb{R}^{b}$, which may include phenomena such as decision-making strategies, herding behavior, but also panic and irrationality [65, 66, 67, 68, 69], where $b\in\mathbb{R}$ is the number of behavioral traits modeled for each pedestrian. Additionally, we introduce an auxiliary variable $\bm{z}_{i}\in\mathbb{R}^{b}$ to denote the rate of change of the behavioral variables $\bm{u}_{i}$. Then, the evolution of the position, velocity and behavior of each pedestrian is modelled by an individualistic/agent-based model, inspired by pseudo-Newtonian/molecular/Brownian dynamics formulations [70, 16], expressed in the general form:
@@
In contrast, at the macroscopic scale, crowd dynamics are typically described using phenomenological continuum models, most often formulated as partial differential equations (PDEs). These models capture the collective behavior of large-scale pedestrian flows and are grounded in the theories of hydrodynamics and continuum mechanics. Well‑known examples include the so‑called Hughes model [17] and multi‑population nonlocal models [18]. However, such modelling approaches rely on several simplifying assumptions in order to derive algebraic closure relations that link individual-level motion with density‑scale dynamics. These assumptions often include, for example, that pedestrian speed depends solely on the local total density, that individuals aim to minimize travel time while avoiding congestion, or that their motion is guided by a potential field directing them toward their destination.
To address the limitations of phenomenological continuum models, several data-driven computational methods have been introduced to construct macroscopic hybrid models in the form of “black-box” or “grey-box” PDE models [50, 51, 49, 52, 53, 23, 42, 43, 21]. These methods are predominantly based on DNNs/NOs, thus suffering from the “curse of dimensionality” that is inherent in their training. To circumvent this challenge, learning Reduced Order Models (ROMs) in well-constructed latent spaces is crucial, since it significantly lowers computational cost while capturing the essential dynamics. These ROMs must also inherently preserve fundamental principles like mass conservation, ensuring physical accuracy in the reconstructed space. Towards this goal, we aim at bridging high-fidelity microscopic/agent-based simulations and/or experimental/real-world data with macroscopic scale modelling, using ROMs that preserve mass conservation, within the framework of learning discrete operators/coarse-timesteppers [35].
The proposed methodology, presented in detail in the following section, is based on our previous efforts to develop what we call next generation Equation-free framework [56, 36, 57, 23, 58]. This framework deals with the “curse of dimensionality” by learning surrogate ML models in appropriate latent spaces emerging from the complex spatio-temporal dynamics of high-dimensional systems. It originates from the celebrated Equation-free approach [35], which enables high-dimensional simulations to perform system-level numerical analysis tasks on an appropriate latent space. This is achieved via the construction of a lifting operator $\mathcal{L}$ and a restriction operator $\mathcal{R}$ bridging the gap between states in a high-dimensional space, say $\mathbf{u}(t)\in\mathbb{R}^{M}$, and states in a latent space, say $\mathbf{y}_{d}(t)\in\mathbb{R}^{d}$, with $d\ll M$, as:
@@
A key distinction between our proposed next generation EF and the traditional EF framework is that, in the latter, one constructs local (numerical) maps and computes the quantities needed for bifurcation analysis on demand via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured usually by a few moment distributions). In contrast, our proposed method learns a “global dynamical model“ in Eq. (5) via machine learning from long-term, high-fidelity simulations spanning the entire phase space, in low-dimensional latent spaces. This global dynamical-systems-based perspective enables the use of both linear (such as POD) and nonlinear manifold-learning techniques (such as Diffusion maps [71, 72]), and techniques for the solution of the pre-image problem, such as geometric harmonics [73, 74] or autoencoders [75, 76], to construct maps between the high-dimensional and the latent space. The methodology is described analytically in the following section.
The key advantage of our next generation EF framework is its ability to circumvent the need for explicit derivation or closure assumptions in PDEs—a process that inherently introduces biases in modelling crowd density dynamic. Instead, it directly learns the “actual” solution operator in Eq. (5)–-the map that advances observed macroscopic distributions (here the density) forward in time. This learned operator, embedded within computationally efficient linear or nonlinear autoregressive ROMs (MVAR or LSTM, respectively), effectively encodes the complex, often analytically intractable, dynamics of the crowd. By focusing on learning how densities evolve (the solution operator) rather than learning governing equations in the form of PDEs, the method circumvents the significant challenges of the “curse of dimensionality” associated with DNNs/NOs training of such PDEs. The resulting ROMs thus act as practical, data-driven surrogates that implicitly contain the action of the unknown PDE’s solution operator, enabling accurate prediction of future crowd states solely from observed density snapshots. Importantly, we demonstrate that for the proposed configuration, the POD method (used for identifying the latent space) conserves explicitly the mass of the reconstructed density field resulting from the ROMs. An explicit mass conservation of the continuum representations is an essential physical constraint for any credible surrogate model of crowd dynamics.

### v2

@@ Abstract
Bridging the microscopic/individual and macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control. Here, we propose a combined manifold and machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations. The proposed framework builds upon our previous works on the development of next-generation Equation-free algorithms for dealing with the “curse of dimensionality” when learning surrogate models for high-dimensional and multiscale systems. The proposed approach is a four-stage one, explicitly conserving the mass of the reconstructed dynamics in the high-dimensional space. In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation. In the second step, based on manifold learning, we construct a map from the macroscopic ambient space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distributions. The third step involves learning reduced-order surrogate models in the latent space using machine learning techniques, particularly Long Short-Term Memory (LSTMs) networks and Multivariate Autoregressive models (MVARs). Finally, we reconstruct the crowd dynamics in the high-dimensional space in terms of macroscopic density profiles. We demonstrate that the POD reconstruction of the density distribution via Singular Value Decomposition conserves the mass. With this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation for the evolution of the density distribution. For our illustrations, we use the Social Force Model to generate data in a corridor with an obstacle, imposing periodic boundary conditions. The numerical results demonstrate high accuracy, robustness, and generalizability, thus allowing for fast and accurate modelling/simulation of crowd dynamics from agent-based simulations. Notably, linear MVAR models surpass nonlinear LSTMs in predictive accuracy, while also offering significantly lower complexity and greater interpretability.
Keywords Crowd Dynamics, Multiscale modelling, Machine Learning, Numerical Analysis, Conservation laws, Reduced Order Models
@@ 1 Introduction
An important open challenge in the modelling of human crowd dynamics lies in bridging across the different modelling scales. Typically, models are formulated at three distinct levels: the microscopic (individual), mesoscopic (kinetic), and macroscopic (hydrodynamic) scales. Microscopic models describe the detailed behavior of individual pedestrians using agent-based approaches, such as the Social Force Model (SFM) [1] and its variants (see e.g. [2, 3]), cellular automata [4, 5, 6], etc. The mesoscopic scale serves as an intermediary between the microscopic and macroscopic approaches [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. At the macroscopic scale, modelling describes the emergent/collective behavior of crowds. For instance, in the seminal work of Hughes [17], the author models crowd dynamics by coupling a conservation law with an eikonal equation to represent how individuals choose paths based on a potential field. Moreover, in [18] the authors proposed nonlocal conservation laws for traffic and crowd dynamics, where the flux depends not only on local density but also on averaged information from the surrounding region. In [11], a framework for macroscopic crowd models was presented based on balance laws and kinetic theory, emphasizing how individual-level interactions give rise to complex collective behavior. More recently, [16] highlights new trends such as multiscale coupling, behavioral heterogeneity, and nonlocal interaction kernels.
@@
A major objective, when high-fidelity microsocpic/individual/agent-based simulators and/or detailed experimental data are available, is to derive mesoscopic or macroscopic models that can then be used for numerical analysis, optimization and control of the emergent/collective dynamics, for instance, in the optimal design of spaces, the reduction of evacuation times, and the control of the flows [11, 9, 19, 20, 21, 22, 23, 24]. In this context, kinetic models serve as an intermediate bridge. They are typically derived analytically using statistical mechanics techniques, starting from microscopic interaction rules. These kinetic formulations, in turn, facilitate the derivation of macroscopic (continuum) models through asymptotic methods [10, 16].
@@
The transition from high-dimensional agent-based representations to macroscopic or hydrodynamic descriptions is commonly achieved through closure techniques. These methods establish relationships between higher-order moments of the microscopic distributions and a limited, smaller set of lower-order moments that capture the observable emergent behavior [25, 26, 27]. However, such closures inherently rely on assumptions that introduce biases both for modelling and analysis tasks. Typically, these assumptions include infinitely large populations, homogeneous agents, and uniform and/or local interaction networks, but also unknown “behavioral” characteristics, which can significantly impact the accuracy and validity of collective-level analyses.
@@
Recently, data-driven computational approaches based on physics-free machine learning (ML) have been employed for modelling, numerical analysis, and control of pedestrian, crowd, and traffic dynamics. In this context, a wide range of ML architectures has been explored, including deep neural networks (DNNs) [28], Long Short-Term Memory (LSTM) networks [29, 30], Generative Adversarial Networks (GANs) [31, 32], Convolutional Neural Networks (CNNs) [33, 34, 35], and reinforcement learning methods [36, 37]. For a comprehensive review of these and related ML approaches, see [38]. These approaches have proven effective in forecasting individual and group trajectories within the framework of time-series modelling and forecasting, without relying on physics-based formulations. However, at the same time, such approaches lack interpretability and, importantly, do not incorporate information of the global spatio-temporal emergent dynamics, such as density fields, that are essential in modelling crowd dynamics.
@@
An alternative data-driven paradigm that bypasses the need to explicitly learn surrogate models is the Equation-free multiscale framework [39]. By appropriately coupling on-demand microscopic simulations with coarse-graining, or “restriction”, and subsequent fine-graining, or “lifting”, this approach builds a coarse time-stepper for the evolution at the macroscopic scale, thus bypassing the need for analytically derived continuum equations [40, 41, 42, 43, 44]. Although the equation-free framework enables powerful multiscale numerical analysis and control, it is fundamentally based on the on-demand construction of local discrete maps rather than on deriving a global macroscopic model of the crowd or collective dynamics (for example, a mass-conserving partial differential equation (PDE)). That reliance on local maps reduces interpretability and physical insight, and it hinders reliable generalization in long-term simulations.
@@
To address the aforementioned challenges, a few recent studies have utilized physics-informed neural networks (PINNs) [45] including physics-informed neural operators (NOs) [46, 47] for solving the inverse problem, i.e., that of constructing hybrid models in the form of “grey-box” PDE models from data [48, 49, 25, 27, 47]. Such an approach, combines ML with underlying knowledge of physical principles, such as conservation laws [50] and closures [51], to create hybrid models that can both predict the spatio-temporal dynamics and ensure physical consistency, even under data scarcity. In the context of crowd dynamics, PINNs have been employed to construct both gray-box hydrodynamic PDEs [52] as well as for discovering the physics/rules governing interactions and motion at the microscopic level [53]. However, while PINNs provide an effective way to incorporate physical constraints and knowledge into ML models, they can also introduce biases due to their reliance on predefined forms of PDEs. Moreover, a key physical property in crowd modeling, mass conservation, is typically not enforced explicitly, but rather imposed as a soft constraint through the loss function [54, 50, 55]. This indirect enforcement can limit the generalizability and physical reliability of PINN-based models, particularly in heterogeneous or unforeseen scenarios where the true dynamics deviate from the assumed PDE forms.
@@
For learning classes of PDEs that incorporate conservation laws–such as those found in crowd dynamics–the use of black-box DNNs and black-box NOs [56, 57, 55, 58, 59, 27] is inherently problematic. These models often lack physical consistency, particularly in preserving conservation laws, which are usually enforced only as soft constraints in the loss function. In addition, their training suffers from low approximation accuracy due to the “curse of dimensionality”, potentially leading to unphysical predictions and numerical instabilities during long-term simulations. Thus, to deal with that problem, RiemannONets were recently presented in [60], which are NOs based on DeepONets to solve Riemann problems encountered in compressible flows, thus enforcing the preservation of conservation laws. Towards solving the inverse problem, we have recently proposed GoRINNs [61], neural networks coupled with Godunov numerical schemes and Riemann solvers for learning hyperbolic PDEs that explicitly respect the conservation laws by construction.
@@
It is important to note, however, that despite the long-standing tradition of modeling crowd dynamics through (hydrodynamic) PDEs, such formulations inherently rely on an “infinite-size” assumption. In realistic settings involving crowds of moderate or finite-size, this assumption often breaks down. Finite-size effects, such as spatial heterogeneity, individual interactions, and boundary influences, can dominate the dynamics, rendering continuum PDEs insufficient for accurate, quantitative predictions of real-world crowd behavior. Here, inspired by the Equation-free multiscale framework [39], and building on our previous recent works, on what we call next-generation Equation-free algorithms based on both manifold learning and machine learning [62, 42, 63, 27, 64, 47], we present a computational framework, to learn the discrete evolution operator of the emergent dynamics in latent spaces from high-fidelity microscopic simulations of crowd dynamics. The proposed framework bridges the microscopic and macroscopic modelling scales, while explicitly preserving the underlying conservation law of mass in the ambient space. The methodology is deployed in four main stages. First, we map discrete microscopic distributions to continuous macroscopic representations (i.e., density fields). Second, we project the macroscopic fields into a low-dimensional latent space using Proper Orthogonal Decomposition (POD). Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM models. These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’ theorem [65] for reconstructing the phase-space structure (here in the latent space). The concept of the approach (i.e., using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [66, 67] (see also results and discussion in [67, 62, 68, 69, 63]). In crowd dynamics, behavioral characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Takens’ theorem, to reconstruct the effective phase space containing these (unobservable) latent dynamics. Finally, we lift the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis. Crucially, this configuration ensures explicit mass conservation—a key physical constraint for any effective continuum description. For our numerical experiments, we employ the Social Force Model (SFM) [70] to simulate pedestrian flow within a rectangular corridor containing an obstacle, under periodic boundary conditions. The performance and generalization ability of the proposed framework are assessed on unseen initial conditions.
@@ 2.1 Problem Statement
We utilize a detailed simulator that models the dynamics of a human crowd, consisting of $N$ pedestrians moving within a domain $\Omega\subset\mathbb{R}^{2}$. The state of each pedestrian $i\in\{1,\ldots,N\}$ is characterized by its position $\mathbf{x}_{i}=\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ and velocity $\mathbf{v}_{i}=\mathbf{v}_{i}(t)=(v_{ix}(t),v_{iy}(t))$, and also a set of “behavioral” variables $\bm{u}_{i}=\bm{u}_{i}(t)\in\mathbb{R}^{b}$, which may include phenomena such as decision-making strategies, herding behavior, but also panic and irrationality [71, 72, 73, 20, 74], where $b\in\mathbb{N}$ is the number of behavioral traits modeled for each pedestrian. Additionally, we introduce an auxiliary variable $\bm{z}_{i}\in\mathbb{R}^{b}$ to denote the rate of change of the behavioral variables $\bm{u}_{i}$. Then, the evolution of the position, velocity and behavior of each pedestrian is modelled by an individualistic/agent-based model, inspired by pseudo-Newtonian/molecular/Brownian dynamics formulations [75, 16], expressed in the general form:
@@
In contrast, at the macroscopic scale, crowd dynamics are typically described using phenomenological continuum models, most often formulated as PDEs. These models capture the collective behavior of large-scale pedestrian flows and are grounded in the theories of hydrodynamics and continuum mechanics. Well‑known examples include the so‑called Hughes model [17] and multi‑population nonlocal models [18]. However, such modelling approaches rely on several simplifying assumptions in order to derive algebraic closure relations that link individual-level motion with density‑scale dynamics. These assumptions often include, for example, that pedestrian speed depends solely on the local total density, that individuals aim to minimize travel time while avoiding congestion, or that their motion is guided by a potential field directing them toward their destination.
To address the limitations of phenomenological continuum models, several data-driven computational methods have been introduced to construct macroscopic hybrid models in the form of “black-box” or “grey-box” PDE models [56, 57, 55, 58, 59, 27, 48, 49, 25]. These methods are predominantly based on DNNs/NOs, thus suffering from the “curse of dimensionality” that is inherent in their training. To circumvent this challenge, learning ROMs in well-constructed latent spaces is crucial, since it significantly lowers computational cost while capturing the essential dynamics. These ROMs must also inherently preserve fundamental principles like mass conservation, ensuring physical accuracy in the reconstructed space. Towards this goal, we aim at bridging high-fidelity microscopic/agent-based simulations and/or experimental/real-world data with macroscopic scale modelling, using ROMs that preserve mass conservation, within the framework of learning discrete operators/coarse-timesteppers [39].
The proposed methodology, presented in detail in the following section, is based on our previous efforts to develop what we call the next generation Equation-free (EF) framework [62, 42, 63, 27, 64]. This framework deals with the “curse of dimensionality” by learning surrogate ML models in appropriate latent spaces emerging from the complex spatio-temporal dynamics of high-dimensional systems. It originates from the Equation-free approach [39], which enables high-dimensional simulations to perform system-level numerical analysis tasks on an appropriate latent space. This is achieved via the construction of a lifting operator $\mathcal{L}$ and a restriction operator $\mathcal{R}$ bridging the gap between states in a high-dimensional space, say $\mathbf{u}(t)\in\mathbb{R}^{M}$, and states in a latent space, say $\mathbf{y}_{d}(t)\in\mathbb{R}^{d}$, with $d\ll M$, as:
@@
A key distinction between our proposed next generation EF and the traditional EF framework is that, in the latter, one constructs local (numerical) maps and computes the quantities needed for bifurcation analysis on demand via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured usually by a few moment distributions). In contrast, our proposed method learns a “global dynamical model” in Eq. (5) via ML from long-term, high-fidelity simulations spanning the entire phase space, in low-dimensional latent spaces. This global dynamical-systems-based perspective enables the use of both linear (such as POD) and nonlinear manifold-learning techniques (such as Diffusion maps [76, 77]), and techniques for the solution of the pre-image problem, such as geometric harmonics [78, 79] or autoencoders [80, 81], to construct maps between the high-dimensional and the latent space. The methodology is described in detail in the following section.
The key advantage of our next generation EF framework is its ability to circumvent the need for explicit derivation or closure assumptions in PDEs—a process that inherently introduces biases in modelling crowd density dynamics. Instead, it directly learns the “actual” evolution operator in Eq. (5)–-the map that advances observed macroscopic distributions (here the density) forward in time. This learned operator, embedded within computationally efficient linear or nonlinear autoregressive ROMs (MVARs or LSTMs, respectively), effectively encodes the complex, often analytically intractable, dynamics of the crowd. By focusing on learning how densities evolve (the evolution operator) rather than learning governing equations in the form of PDEs, the method circumvents the significant challenges of the “curse of dimensionality” associated with DNNs/NOs training of such PDEs. The resulting ROMs thus act as practical, data-driven surrogates that implicitly contain the action of the unknown PDE’s solution operator, enabling accurate prediction of future crowd states solely from observed density snapshots. Importantly, we demonstrate that for the proposed configuration, the POD approach (used for identifying the latent space) conserves explicitly the mass of the reconstructed density field resulting from the ROMs. An explicit mass conservation of the continuum representations is an essential physical constraint for any credible surrogate model of crowd dynamics.